Word problem and having a hard time setting this up. I want to write it as, 40=x+(5+y) what am I doing and what should I be doing?
THanks
Admission to the zoo costs $5 more for an adult than for a child. Admission for two adults and two children costs $40. Find the cost of each adults admission and each childs admission.

Answer by ankor@dixie-net.com(22740)

Admission to the zoo costs $5 more for an adult than for a child.
Admission for two adults and two children costs $40.
Find the cost of each adults admission and each childs admission.
:
Let c = child's price of admission
then
(c+5) = adults price
:
2c + 2(c+5) = 40
2c + 2c + 10 = 40
4c = 40 - 10
4c = 30
c = 30/4
c = $7.50 for 1 child
then
7.50 + 5 = $12.50 for an adult